OpenStudy (firejay5):

Determine the slope of a line perpendicular to the line with equation: 6x - 4y = 30. Show your work and/or explain how you did it. Medal will be rewarded for your patronage and accuracy & correctness of answer. Thanks! :DDD

OpenStudy (texaschic101):

do you know how to find the slope in the equation 6x - 4y = 30 ?

OpenStudy (texaschic101):

put the equation in y = mx + b form, and the slope of the line is in the m position. Once you find the slope, we will work on finding the slope of the perpendicular line.

OpenStudy (firejay5):

Well I just switched the equation into slope-intercept form: y = mx + b

OpenStudy (texaschic101):

what did you get ?

OpenStudy (firejay5):

y = -3/2 - 15/2

OpenStudy (anonymous):

Solve for Y. The x coefficient will be the slope of the line. -4y = -6x + 30 (subtract 6x from both sides) y = 3/2x - 15/2 (Divide both sides by -4) Slope is 3/2

OpenStudy (texaschic101):

6x - 4y = 30 -4y = -6x + 30 y = -6/-4y + 30/-4 y = 3/2 - 15/2 the slope of the equation you were given is 3/2 for a perpendicular line, you have to take the negative reciprocal of that slope. To do that, you " flip " the fraction and change the sign. Therefore, the negative reciprocal of 3/2 = -2/3. so the slope of the perpendicular line you are looking for is -2/3. Do you understand ?

OpenStudy (firejay5):

\[6x - 4y = 30 --> \frac{ -4y }{ -4 } = \frac{ 6x }{ -4 } -\frac{ 30 }{ 4 }\]

OpenStudy (texaschic101):

6x - 4y = 30 (subtract 6x from both sides) 6x - 6x - 4y = -6x + 30 (combine like terms) -4y = -6x + 30 divide by -4 to get y by itself y = (-6/-4)y + (30/-4) y = 3/2y - 15/2 slope is 3/2 negative reciprocal of 3/2 = - 2/3 (your perpendicular slope)

OpenStudy (texaschic101):

oops...typo ...y = 3/2x - 15/2

OpenStudy (firejay5):

or just change it into slope - intercept form: y = mx + b

OpenStudy (texaschic101):

yes...y = mx + b form would be y = 3/2x - 15/2 and 3/2 is the slope of this equation. perpendicular slope would be negative reciprocal of 3/2, which is -2/3

OpenStudy (firejay5):

I got -3/2 - 15/2

OpenStudy (texaschic101):

it is not -3/2...it is 3/2. I worked the problem out above....

OpenStudy (firejay5):

It is, because -4y/4 = y; 6x/-4 = -3/2, 30/-4 = -15/2 or -7.5

OpenStudy (texaschic101):

let me show you again.... 6x - 4y = 30 subtract 6x from both sides 6x - 6x - 4y = -6x + 30 leaving you with -4y = -6x + 30 now we divide by -4 -4y/-4 = -6x/-4 + 30/-4 y = 3/2x - 15 you are messing up on dividing -6/-4, two negatives divided by each other = a positive.

OpenStudy (texaschic101):

understand ?

OpenStudy (firejay5):

no, how about solving in by entering 6x - 4y = 30 into the slope intercept form: y = mx + b and solve it that way

OpenStudy (texaschic101):

that is what we just did.... y = 3/2x - 15 is slope intercept form. and to find perpendicular slope, you take the negative reciprocal of slope 3/2 which is -2/3

OpenStudy (firejay5):

no you didn't; I'll get you started: -4y = 6x + 30, go from there

OpenStudy (texaschic101):

no...... you have to subtract 6x from both sides to get -4y by itself. therefore, it becomes -4y = -6x + 30

OpenStudy (texaschic101):

@RadEn ...can you explain this...he does not seem to understand me

OpenStudy (firejay5):

It should be -15/4 not -15

OpenStudy (texaschic101):

@ybarrap ...can you explain this...he is not understand me

OpenStudy (firejay5):

I understand it, you made an error -15 is wrong; it's -15/2

OpenStudy (texaschic101):

oops....-15/2.....however, that does not matter in this problem

OpenStudy (texaschic101):

we are just dealing with the slope, not the y intercept

OpenStudy (firejay5):

I completely understand it, I just had to point that out

OpenStudy (texaschic101):

do you understand how the slope is 3/2 and not -3/2 ?

OpenStudy (ybarrap):

Great job guys!!

OpenStudy (firejay5):

so the answer is -2/3, because the slope of the equation was 3/2x - 15/2

OpenStudy (texaschic101):

good deal :) yes...the perpendicular slopeis -2/3 Just remember, to find perpendicular, you take the negative reciprocal of the slope given

OpenStudy (texaschic101):

thanks ybarrap :)